How many moles of PCl5 can be produced from 26.0 g of P4 (and excess Cl2)

P4(s)+10Cl2(g)→4PCl5(g)
Answer #1

molar mass of P4    = 4*31   = 124g/mole

mass of P4             = 26g

no of moles of P4    = mass of P4/gram molar mass of P4

                                 = 26/124

                                  = 0.2097 moles

P4(s)+10Cl2(g)→4PCl5(g)

1 mole of P4 react with excess of Cl2 to gives 4 moles of PCl5

0.2097 moles of P4 react with excess of Cl2 to gives = 4*0.2097/1   = 0.839moles of PCl5 >>>>>answer